OpenStudy (shayaan_mustafa):

Differential Calculus: The minute hand of a certain clock is 4 in long. Starting from the moment when the hand is pointing straight up, how fast is the area of the sector that is swept out by the hand increasing at any instant during the next revolution of the hand?

OpenStudy (shayaan_mustafa):



OpenStudy (shayaan_mustafa):

Now help me from here. I have made diagram. Area of the sector is, \(\Large\frac{\theta}{2}r^2\)

OpenStudy (shayaan_mustafa):

Kindly tell me what is problem is trying to say. actually I am not getting the question.

OpenStudy (anonymous):

Hint: Divide the entire clock area into 60 equal sectors or units. Then area increasing at the rate of one unit/sector per second. Clear?

OpenStudy (anonymous):

One unit or sector = ( Pi r ^2 )/60

OpenStudy (shayaan_mustafa):

\[\Large A=\frac{\theta}{2}r^2\] \[\Large \frac{dA}{dt}=\frac{1}{2}[r^2.\frac{d\theta}{dt}+2r.\theta.\frac{dr}{dt}]\] Is it correct?

OpenStudy (shayaan_mustafa):

@Champs Why divide into 60 equal sectors?

OpenStudy (turingtest):

mustafa I think what you have is correct, but note that dr/dtheta=0

OpenStudy (shayaan_mustafa):

OK OK unit/sector. That is why I should divide it? Right?

OpenStudy (turingtest):

dr/dt=0*

OpenStudy (shayaan_mustafa):

How dr/dt= 0 ?

OpenStudy (turingtest):

r is the length of the hand of the clock it's not getting any longer or shorter , is it?

OpenStudy (shayaan_mustafa):

actually I am not getting this part of the question, below; "how fast is the area of the sector that is swept out by the hand increasing at any instant during the next revolution of the hand?"

OpenStudy (shayaan_mustafa):

Yes you are right. Minute hand's length is fixed i.e. 4 inch.

OpenStudy (turingtest):

the "during the next revolution" part is poor english I don't know who translated that part but it is confusing a "revolution" is all the way around...

OpenStudy (turingtest):

right, so your answer makes sense to me up to\[\frac{dA}{dt}=\frac12r^2\frac{d\theta}{dt}\]and what is \(\large \frac{d\theta}{dt}\)

OpenStudy (shayaan_mustafa):

during next revolution means, what I understood is, firstly one complete rotation and then it is taking about 2nd revolution, isn't it? 1 revolution = 2\(\Large \pi\)

OpenStudy (turingtest):

yes, but the rate of increase in the area swept out by the hand is a constant because 1/2, r^2, and d(theta)/dt are all constants, so what is the point of phrasing the question like that?

OpenStudy (shayaan_mustafa):

We have given \(\theta\) i.e. 1 revolution=2\(\Large \pi\) But we have not given \(\Large \frac{d\theta}{dt}\)

OpenStudy (turingtest):

but you can figure it out...

OpenStudy (turingtest):

it's a clock, and as far as I know \[\frac{d\theta}{dt}\]is the same for all the working clocks in the world!

OpenStudy (shayaan_mustafa):

So what it should be? Because I am not getting. If I differentiate it with respect to time\(\large2\pi\). Then it is a constant so it will go 0.

OpenStudy (turingtest):

remember what \[\frac{d\theta}{dt}\]represents; the rate change in the angle i.e. (angle/time) the angle for a revolution of a clock is \(2\pi\) radians and the time it takes to make one revolution is 60seconds. Since the hand is moving at a constant rate, the average angular change is the same as the instantaneous one (the derivative).

OpenStudy (shayaan_mustafa):

Why 60sec? It is a minute hand. So it will 3600sec for one revolution.. Isn't it?

OpenStudy (turingtest):

ah good point still it's a constant rate so my argument persists with different numbers\[\frac{d\theta}{dt}=\frac{\Delta\theta}{\Delta t}=\frac{2\pi}{3600}=\frac\pi{1800}\]if the time is in seconds

OpenStudy (shayaan_mustafa):

OK great thinking. Let me solve what we have until do. Then I post the result. Wait then. please be online.

OpenStudy (turingtest):

sure :)

OpenStudy (shayaan_mustafa):

\[\frac{dA}{dt}=\frac{1}{2}r^2\frac{d\theta}{dt}\] \[\frac{dA}{dt}=\frac{1}{2}(16)(\frac{\pi}{1800})\] \[\frac{dA}{dt}=0.014in^2/min\] Is this right?

OpenStudy (turingtest):

that's what I got or about 0.84in^2/min you never specified the units you wanted

OpenStudy (turingtest):

your time is in seconds I think

OpenStudy (shayaan_mustafa):

theta is in radians.

OpenStudy (shayaan_mustafa):

no. time in minute.

OpenStudy (shayaan_mustafa):

see i have shown unit. it is in^2/min.

OpenStudy (turingtest):

then your answer is wrong

OpenStudy (shayaan_mustafa):

Why? should theta be in degree? or time in sec?

OpenStudy (shayaan_mustafa):

I have made calculation above. Spot the mistake please.

OpenStudy (turingtest):

you have time in seconds as \(you\) were actually the one to poinn out to me, lol

OpenStudy (turingtest):

point*

OpenStudy (turingtest):

if you want the units in radians per minute then you need to calculate \[\frac{d\theta}{dt}\]in minutes and how many minutes per revolution of the minute hand?

OpenStudy (shayaan_mustafa):

Wait wait. I used \(\large \frac{d\theta}{dt}\)=\(\large \frac{\pi}{1800}\). So I think it is radian per sec. Right?

OpenStudy (turingtest):

exactly

OpenStudy (turingtest):

that's 'cuz we did divided 2pi by 3600 since there are 3600 seconds in a revolution if we want it in minute we have to divide by the number of minutes in a revolution, which is...?

OpenStudy (shayaan_mustafa):

I am getting confused now.

OpenStudy (shayaan_mustafa):

I have made calculation. Did you check them?

OpenStudy (turingtest):

if you want radians per second\[\frac{d\theta}{dt}=\frac{\Delta\theta}{\Delta t}=\frac{2\pi\text{ rad}}{3600\sec}=\frac\pi{1800}\frac{\text{rad}}{\sec}\]if you want radians per minute:\[\frac{d\theta}{dt}=\frac{\Delta\theta}{\Delta t}=\frac{2\pi\text{ rad}}{60\min}=\frac\pi{30}\frac{\text{rad}}{\min}\]you just have to be conscious of your units!

OpenStudy (turingtest):

...because it is 3600seconds=60minutes for the minute hand to go around once

OpenStudy (turingtest):

and yes I have checked your answers and that is your only mistake

OpenStudy (shayaan_mustafa):

OK.. That is what I don't know, that I should consider units too. Thank you. It is done now.

OpenStudy (turingtest):

welcome :)